Question regarding stock market pricing, I need help with. Thank you in advance!

Given that the price a stock is bought for is $100. Based on the one-period valuation model of stock prices, if the stock is sold a year later at the price of $125 after receiving a dividend of $2, then the required rate of return on equity investments is __%. (Round your response to the nearest one decimal percentage)

Now, suppose that the price of the stock above was bought instead for $115. Then, required rate of return on equity investments then ___ (the possible answers for this got cutoff so I'm not sure what the options for this blank are, any predictions would be much appreciated!)
Answer #1

Required rate of return in the first case:

Return on sale = $25

Dividend= $2

Hence return on Investment = Total return/ Investment value*100 = 27/100*100= 27%pa

Required rate of return in the first case:

Return on sale = $ 10

Dividend = $2

Hence return on Investment = Total return/ Investment value*100 = 12/100*100= 12%pa

Alternative:

If it is considered that the dividend is growing at a constant rate of 8%, then

Required rate of return = (Dividend/ Current stock value)+Growth = (2/100)+8%=10%

In the second case Required rate of return = (2/115) + 8%= 9.74%